# A Survey on Error-bounded Lossy Compression for Scientific Datasets

### v2

@@ 1. Introduction
In this paper we present a comprehensive survey to provide a thorough understanding of the error-controlled lossy compression techniques and how they are used in different parallel and distributed use cases. The following topics are considered.
@@
To the best of our knowledge, this is the most comprehensive summary of the lossy compression modules/techniques used by existing error-bounded lossy compressors (up to the year 2024), and the most comprehensive survey for the emerging state-of-the-art error-bounded lossy compressors.
@@ 3. Compression Model Taxonomy
Figure 2. Compression Pipeline with Various Models: each highlighted box represents the key step of the corresponding model. All the compression techniques shown here will be detailed in Section 4.
In Figure 1, each model has its pros and cons with respect to the time/space complexity and reconstruction quality. The models on the left generally tend to have lower time complexity yet lower data reconstruction quality than do the models on the right. [^2] Note that the performance and quality rule demonstrated in Figure 1 holds in general cases. The real performance/quality also depends on the specific design and implementation. In practice, each reduction model is a fundamental technique that can be combined with other models or techniques to generate a specific data compressor, which will be detailed in Section 5.
@@ 4. Modular Lossy Compression Techniques
In this section we describe the key lossy compression modules/techniques often used in many modern state-of-the-art lossy compressors. Each technique listed here serves as just a module or step in a compression pipeline (i.e., a compressor). That is, each technique often needs to be combined with one or more other techniques to compose an error-bounded lossy compression pipeline in order to obtain a high compression ratio, forming a many-to-many relationship between compression-pipeline and compression-technique. Figure 3 presents the usage distribution of compression techniques in the 46 lossy compressors we collected.
@@ 5. General-Purpose Lossy Compressors for Scientific Data
In this section, we describe general-purpose lossy compressors, each of which projects a specific compression pipeline composed of a series of steps or techniques/modules. The selection of lossy compressorsin our study is based on the following criteria.
  • These compressors are designed for scientific datasets generated by HPC applications or advanced X-ray facilities such as APS (Fornek, 2017).
  • These compressors provide certain compression error controls, in the form of error bound (Di and Cappello, 2016; Tao et al., 2017c), bit-truncation precision (Lindstrom et al., 2017; Lindstrom, 2014), or control of impact to the posthoc analysis.
@@
Table 6. Summary of General-purpose Lossy Compressors
Lossy Compressors Year Model Compress. Pipeline Device Type QoI Performance Compression Ratio FPZIP (Lindstrom et al., 2017) 2006 Pred. PDP+BPC CPU S N *** (Di and Cappello, 2016) ** (Liang et al., 2018b; Di and Cappello, 2016) ISABELA (Lakshminarasimhan et al., 2011, 2013) 2013 Deci. Sorting+DS+Bspline CPU S N * (Di et al., 2018) * (Di et al., 2018) ZFP (Lindstrom, 2014; cuZFP, 2020) 2014 Trans. DT+OWT+BPC CPU/GPU S N **** (Tao et al., 2017c; Liang et al., 2018b) ****(Tao et al., 2017c; Liang et al., 2018b) NUMARCK (Chen et al., 2014) 2014 Pred. PDP+QT CPU S N - ** (Di and Cappello, 2016) SSEM (Sasaki et al., 2015) 2015 Trans. DWT+QT+FTR+LE CPU S N - ** (Di and Cappello, 2016) SZ0.1 (Di and Cappello, 2016) 2016 Pred. PDP+LE CPU S N *** (Di and Cappello, 2016) *** (Di and Cappello, 2016) Bitgrooming (Zender, 2016) 2016 BitM. BPC+LE CPU S N - * (Underwood et al., 2022) SZ1.4 (Tao et al., 2017c) 2017 Pred. PDP+QT+LE CPU S N *** (Tao et al., 2017c; Liang et al., 2018b) *** (Tao et al., 2017c; Liang et al., 2018b) SZ2 (Liang et al., 2018b) 2018 Pred. PDP+QT+LE CPU S N **** (Liang et al., 2018b; Zhao et al., 2021) **** (Liang et al., 2018b; Zhao et al., 2021) MGARD (Ainsworth et al., 2018, 2020) 2018 Pred. PDP+QT+LE CPU U&S Y ** (Liang et al., 2021b) *** (Liang et al., 2021b) SpaioTmpDeci. (Liang et al., [n. d.]) 2018 Decim. DS CPU S N **** (Liang et al., [n. d.]) * (Liang et al., [n. d.]) Digitrounding (dig, [n. d.]) 2019 BitM. BPC+LE CPU S N - * (Underwood et al., 2022) DCTZ (Zhang et al., 2019b) 2019 Trans. OWT+QT+LE CPU S N *** (Zhang et al., 2019b) *** (Zhang et al., 2019b) GhostSZ (Xiong et al., 2019) 2019 Pred. PDP+QT+LE FPGA S N *** (Xiong et al., 2019) **** (Xiong et al., 2019) TTHRESH (Ballester-Ripoll et al., 2019) 2019 HOSVD SVD+BPC+LE CPU S N * (Ballester-Ripoll et al., 2019; Liu et al., 2024; Baert and Vannieuwenhoven, 2023) ***** (Baert and Vannieuwenhoven, 2023) ZFP-V (Sun and Jun, 2019) 2019 Trans. DT+OWT+BPC+LE FPGA S N **** (Sun and Jun, 2019) **** (Sun and Jun, 2019) SZauto (Zhao et al., 2020b) 2020 Pred. PDP+QT+LE CPU S N *** (Zhao et al., 2020b) **** (Zhao et al., 2020b) waveSZ (Tian and et al., 2020) 2020 Pred. QT+PDP+LE FPGA S N *** (Tian and et al., 2020) **** (Tian and et al., 2020) TuckerMPI (Ballard et al., 2020) 2020 HOSVD SVD CPU S N ** (Baert and Vannieuwenhoven, 2023) ** (Baert and Vannieuwenhoven, 2023) zMesh(Luo et al., 2021) 2021 Pred. DT CPU U N *** (Luo et al., 2021) **** (Luo et al., 2021) TEZip (Roy and et al., 2021) 2021 DeepL. DNN+LE CPU S N **** (Roy and et al., 2021) **** (Roy and et al., 2021) cuSZ (Tian et al., 2020; Tian and et al., 2021) 2021 Pred. QT+PDP+LE GPU S N ** (Huang et al., 2023a) *** (Huang et al., 2023a, 2024a; Liu and et al., 2024) MGARD+ (Liang et al., 2021b) 2021 Pred. PDP+QT+LE CPU S N *** (Liang et al., 2021b) **** (Liang et al., 2021b) SZ3 (Zhao et al., 2021; Liang et al., 2021a) 2021 Pred. PDP+QT+LE CPU S N *** (Zhao et al., 2021) **** (Zhao et al., 2021; Liu et al., 2024, 2023c) AESZ (Liu et al., 2021a, d) 2021 DeepL. DNN+PDP+QT+LE CPU S N ** (Liu et al., 2021a) **** (Liu et al., 2021a) CAE (Liu et al., 2021d) 2021 DeepL. DNN+PDP+QT+LE CPU S N ** (Liu et al., 2021d) **** (Liu et al., 2021d) MGARD-GPU(Chen et al., 2021; Gong et al., 2023) 2021 Pred. PDP+QT+LE GPU S N *** (Gong et al., 2023) *** (Gong et al., 2023) DE-ZFP(Habboush et al., 2022) 2022 Trans. DT+OWT+BPC+LE FPGA S N **** (Habboush et al., 2022) **** (Habboush et al., 2022) QoZ (Liu et al., 2022b) 2022 Pred. PDP+QT+LE CPU S N *** (Liu et al., 2022b, 2024) **** (Liu et al., 2022b, 2024) SZx (Yu et al., 2022) 2022 BitM. FTR+BPC CPU S N ***** (Yu et al., 2022) ** (Yu et al., 2022) cuSZx (Yu et al., 2022) 2022 BitM. FTR+BPC GPU S N ** (Huang et al., 2023a) ** (Huang et al., 2023a) Jiao et al. (Jiao et al., 2022) 2022 Pred. DT+PDP+QT+LE CPU S Y ** (Jiao et al., 2022) *** (Jiao et al., 2022) MGARD-Lambda (mga, [n. d.]) 2022 Pred. PDP+QT+LE CPU S Y - - ATC (Baert and Vannieuwenhoven, 2023) 2023 HOSVD SVD+QT+BPC+LE CPU S N ** (Baert and Vannieuwenhoven, 2023) ***** (Baert and Vannieuwenhoven, 2023) ZFP-X (Lu et al., 2023) 2023 Trans. DT+OWT+BPC+LE CPU S N **** (Lu et al., 2023) **** (Lu et al., 2023) SPERR (Li et al., 2023) 2023 Trans. DWT+QT+LE CPU S N ** (Li et al., 2023; Liu et al., 2023c) ***** (Li et al., 2023; Liu et al., 2023c) HH-NN (Huang and Hoefler, 2023) 2023 DeepL. DS+DNN+QT GPU S N * (Huang and Hoefler, 2023) ***** (Huang and Hoefler, 2023) FAZ (Liu et al., 2023c) 2023 Pred. DWT+PDP+QT+LE CPU S N ** (Liu et al., 2023c) ***** (Liu et al., 2023c) FZ-GPU (Zhang et al., 2023) 2023 Pred. QT+PDP+LE GPU S N **** (Zhang et al., 2023) ** (Zhang et al., 2023) cuSZp (Huang et al., 2023a, 2024a) 2023 Pred. QT+PDP+LE GPU S N ***** (Huang et al., 2023a, 2024a) ** (Huang et al., 2023a, 2024a) AMR-Comp (Wang et al., 2022, 2024b, 2023b, 2023a, 2024a) 2023 Pred. PDP+QT+LE CPU U N *** (Wang et al., 2024b) **** (Wang et al., 2024b) SRNN-SZ (Liu et al., 2023b) 2023 Pred. DNN+PDP+QT+LE CPU S N * (Liu et al., 2023b) ***** (Liu et al., 2023b) SZ_ADT (Lu and et al., 2023) 2023 Pred. PDP+QT+LE CPU S N *** (Lu and et al., 2023) **** (Lu and et al., 2023) HPEZ(QOZ2) (Liu et al., 2024) 2024 Pred. PDP+QT+LE CPU S N *** (Liu et al., 2024) **** (Liu et al., 2024) HAE (Le et al., 2024) 2024 DeepL. DT+DNN+QT+FTR GPU S N - **** (Le et al., 2024) SZp (szp, [n. d.]; Huang and et al., 2024a) 2024 Pred. QT+PDP+LE GPU S N ***** (Huang and et al., 2024a) ** (Huang and et al., 2024a)
@@
SZ-family compressors. SZ (Di and Cappello, 2016; Tao et al., 2017c; Zhao et al., 2021; Liang et al., 2021a; Di and Cappello, 2018) is a prediction-based error-bounded lossy compressor. In fact, it is not only a compression library/software but also a flexible composable framework allowing users to customize specific compression pipelines according to their datasets or use cases. SZ’s compression pipeline generally is composed of four stages: pointwise data prediction, quantization, variable-length encoding and lossless encoding. For different domain datasets and use cases, the SZ developers have developed many predictors, including Lorenzo, linear regression, dynamic spline interpolation (Zhao et al., 2021), and also using ZFP transform as a predictor (Liang et al., 2019a). In addition to the data prediction stage, SZ developers explored the possibilities of improving compression capability by other stages, for example, different quantization methods such that the users can set various error bounds for multiple value ranges in one dataset (Liu et al., 2021b; Liu and et al., 2021; Liu et al., 2022a). SZ adopts Huffman+Zstd (Di and Cappello, 2016; Tao et al., 2017c; Liang et al., 2018b) to compress the quantization bins because this method projects the best trade-off between the compression ratio and compression speed (Liu et al., 2021c). SZ also supports pointwise relative error bound compression (Liang et al., 2018a), by leveraging a preprocessing step to transform the original data domain to the logarithm domain and then executing absolute error-bounded compression on top of it. To accelerate the throughput, an efficient fusion of logarithm was developed in (Zou et al., 2019, 2020).
@@
Based on SZ3 framework, the SZ team has developed many variants to adapt to diverse user-demands and use-cases. For example, QoZ (quality-oriented compressor) (Liu et al., 2022b) supports compression autotuning according to user-specified quality metric targets. QoZ2.0/HPDZ (Liu et al., 2024) brings several major updates to its data prediction design, including multidimensional interpolation, interpolation re-ordering, dimension autofolding, and blockwise interpolation tuning, which can achieve 50% to 300% compression ratio improvement under the same quality metric (such as PSNR). FAZ (Liu et al., 2023c) combines SZ3 model with wavelet transform and can also adaptively leverage the best-fit compression techniques for each separate input and autodetermine their corresponding parameters. FAZ features extremely high compression ratios over SZ3 ( at a cost of compression throughput): it can get 3-4$\times$ compression ratio as high as that of QoZ on the SEGSalt dataset as shown in (Liu et al., 2023c). In addition, the SZ team also explored the possibility of leveraging DNN to improve compression ratios under the SZ3 framework: e.g., AE-SZ (Liu et al., 2021a) leverages AutoEncoder and SRNN-SZ (Liu et al., 2023b) adopts super-resolution technique. In addition, the SZ team also developed several ultra-fast versions for both CPU and GPU devices. SZx (Yu et al., 2022), for example, composes only lightweight operations, such as bitwise operations, additions, and subtractions, so that it can get a very high compression/decompression speed (about 4$\times$ as fast as SZ2 and ZFP in general). CuSZp (Huang et al., 2023a, 2024a) can get a very high end-to-end performance on GPU for both compression and decompression (200-400GB/s on A100 in general), still maintaining a relatively high compression ratio.
@@
  • Lossless compression: compressing the shuffled bytes with a lossless compressor such as Deflate (Gzip (Deutsch, 1996)) or Zstd (Collet, 2015).

### v3

@@ 1. Introduction
In this paper we present a comprehensive survey to provide a thorough understanding of the error-controlled lossy compression techniques especially for scientific datasets and how they are used in different parallel and distributed use cases.
We identify the scientific datasets targeted here as the datasets generally represented in the form of floating-point or integer values. These datasets could be generated by scientific applications or advanced instruments (such as light-source), and they could be stored in memory for the reuse in the simulation or stored/transferred separately for posthoc analysis. Moreover, these datasets can be structured (e.g., 1D or multi-dimensional grid data) or unstructured (e.g., triangular meshes consisting of nodes, edges, and faces). Specific multi-media datasets (such as image-format or wave-format) are excluded in this survey.
The selection of lossy compressors and research works is based on the following criteria.
  • These compressors are designed for scientific datasets (represented in the above-mentioned formats) generated by HPC applications or advanced X-ray facilities such as APS (Fornek, 2017).
  • These compressors provide certain error control, in the form of error bound (Di and Cappello, 2016; Tao et al., 2017c), bit-truncation precision (Lindstrom et al., 2017; Lindstrom, 2014), or control of impact on posthoc analysis. For example, lossy compression methods such as JPEG and JPEG2000 have been used in certain applications (Zabala et al., 2006), but they are excluded from this survey because they lack explicit mechanisms for controlling or bounding compression errors.
The following topics are considered.
@@
To the best of our knowledge, this is the most comprehensive summary of the compression modules/techniques used by existing error-bounded lossy compressors (up to the year 2024), and the most comprehensive survey for the emerging state-of-the-art error-bounded lossy compressors. Moreover, we highlighted the key takeaways throughout the survey in both italic and bold font to make important insights easier to follow.
@@ 3. Compression Model Taxonomy
Figure 2. Compression Pipeline with Various Models: each highlighted box represents a corresponding model. All the compression techniques shown here will be detailed in Section 4.
Takeaway 1: each model has its pros and cons with respect to the time/space complexity and reconstruction quality. The models on the left generally tend to have lower time complexity yet lower data reconstruction quality than do the models on the right. [^2] Note that the performance and quality rule demonstrated in Figure 1 holds in general cases. The real performance/quality also depends on the specific design and implementation. In practice, each reduction model is a fundamental technique that can be combined with other models or techniques to generate a specific data compressor, which will be detailed in Section 5.
@@ 4. Modular Lossy Compression Techniques
In this section, we describe the key lossy compression modules/techniques often used in many modern state-of-the-art lossy compressors. Each technique listed here serves as just a module or step in a compression pipeline (i.e., a compressor).
Takeaway 2: each compression technique often needs to be combined with one or more other techniques to compose an error-bounded lossy compression pipeline in order to obtain a high compression ratio, forming a many-to-many relationship between compression-pipeline and compression-technique. Figure 3 presents the usage distribution of compression techniques in the 46 lossy compressors we collected.
@@ 5. General-Purpose Lossy Compressors for Scientific Data
Table 6. Summary of General-purpose Lossy Compressors
Lossy Compressors Year Model Compress. Pipeline Device Type QoI Performance Compression Ratio FPZIP (Lindstrom et al., 2017) 2006 Pred. PDP+BPC CPU S N *** (Di and Cappello, 2016) ** (Liang et al., 2018b; Di and Cappello, 2016) ISABELA (Lakshminarasimhan et al., 2011, 2013) 2013 Deci. Sorting+DS+Bspline CPU S N * (Di et al., 2018) * (Di et al., 2018) ZFP (Lindstrom, 2014; cuZFP, 2020) 2014 Trans. DT+OWT+BPC CPU/GPU S N **** (Tao et al., 2017c; Liang et al., 2018b) ****(Tao et al., 2017c; Liang et al., 2018b) NUMARCK (Chen et al., 2014) 2014 Pred. PDP+QT CPU S N - ** (Di and Cappello, 2016) SSEM (Sasaki et al., 2015) 2015 Trans. DWT+QT+FTR+LE CPU S N - ** (Di and Cappello, 2016) SZ0.1 (Di and Cappello, 2016) 2016 Pred. PDP+LE CPU S N *** (Di and Cappello, 2016) *** (Di and Cappello, 2016) Bitgrooming (Zender, 2016) 2016 BitM. BPC+LE CPU S N - * (Underwood et al., 2022) SZ1.4 (Tao et al., 2017c) 2017 Pred. PDP+QT+LE CPU S N *** (Tao et al., 2017c; Liang et al., 2018b) *** (Tao et al., 2017c; Liang et al., 2018b) SZ2 (Liang et al., 2018b) 2018 Pred. PDP+QT+LE CPU S N **** (Liang et al., 2018b; Zhao et al., 2021) **** (Liang et al., 2018b; Zhao et al., 2021) MGARD (Ainsworth et al., 2018, 2020) 2018 Pred. PDP+QT+LE CPU U&S Y ** (Liang et al., 2021b) *** (Liang et al., 2021b) SpaioTmpDeci. (Liang et al., [n. d.]) 2018 Decim. DS CPU S N **** (Liang et al., [n. d.]) * (Liang et al., [n. d.]) Digitrounding (dig, [n. d.]) 2019 BitM. BPC+LE CPU S N - * (Underwood et al., 2022) DCTZ (Zhang et al., 2019b) 2019 Trans. OWT+QT+LE CPU S N *** (Zhang et al., 2019b) *** (Zhang et al., 2019b) GhostSZ (Xiong et al., 2019) 2019 Pred. PDP+QT+LE FPGA S N *** (Xiong et al., 2019) **** (Xiong et al., 2019) TTHRESH (Ballester-Ripoll et al., 2019) 2019 HOSVD SVD+BPC+LE CPU S N * (Ballester-Ripoll et al., 2019; Liu et al., 2024; Baert and Vannieuwenhoven, 2023) ***** (Baert and Vannieuwenhoven, 2023) ZFP-V (Sun and Jun, 2019) 2019 Trans. DT+OWT+BPC+LE FPGA S N **** (Sun and Jun, 2019) **** (Sun and Jun, 2019) SZauto (Zhao et al., 2020b) 2020 Pred. PDP+QT+LE CPU S N *** (Zhao et al., 2020b) **** (Zhao et al., 2020b) waveSZ (Tian and et al., 2020) 2020 Pred. QT+PDP+LE FPGA S N *** (Tian and et al., 2020) **** (Tian and et al., 2020) TuckerMPI (Ballard et al., 2020) 2020 HOSVD SVD CPU S N ** (Baert and Vannieuwenhoven, 2023) ** (Baert and Vannieuwenhoven, 2023) zMesh(Luo et al., 2021) 2021 Pred. DT CPU U N *** (Luo et al., 2021) **** (Luo et al., 2021) TEZip (Roy and et al., 2021) 2021 DeepL. DNN+LE CPU S N **** (Roy and et al., 2021) **** (Roy and et al., 2021) cuSZ (Tian et al., 2020; Tian and et al., 2021) 2021 Pred. QT+PDP+LE GPU S N ** (Huang et al., 2023a) *** (Huang et al., 2023a, 2024a; Liu and et al., 2024) MGARD+ (Liang et al., 2021b) 2021 Pred. PDP+QT+LE CPU S N *** (Liang et al., 2021b) **** (Liang et al., 2021b) SZ3 (Zhao et al., 2021; Liang et al., 2021a) 2021 Pred. PDP+QT+LE CPU S N *** (Zhao et al., 2021) **** (Zhao et al., 2021; Liu et al., 2024, 2023c) AESZ (Liu et al., 2021a, d) 2021 DeepL. DNN+PDP+QT+LE CPU S N ** (Liu et al., 2021a) **** (Liu et al., 2021a) CAE (Liu et al., 2021d) 2021 DeepL. DNN+PDP+QT+LE CPU S N ** (Liu et al., 2021d) **** (Liu et al., 2021d) MGARD-GPU(Chen et al., 2021b; Gong et al., 2023) 2021 Pred. PDP+QT+LE GPU S N *** (Gong et al., 2023) *** (Gong et al., 2023) DE-ZFP(Habboush et al., 2022) 2022 Trans. DT+OWT+BPC+LE FPGA S N **** (Habboush et al., 2022) **** (Habboush et al., 2022) QoZ (Liu et al., 2022b) 2022 Pred. PDP+QT+LE CPU S N *** (Liu et al., 2022b, 2024) **** (Liu et al., 2022b, 2024) SZx (Yu et al., 2022) 2022 BitM. FTR+BPC CPU S N ***** (Yu et al., 2022) ** (Yu et al., 2022) cuSZx (Yu et al., 2022) 2022 BitM. FTR+BPC GPU S N *** (Huang et al., 2023a) ** (Huang et al., 2023a) Jiao et al. (Jiao et al., 2022) 2022 Pred. DT+PDP+QT+LE CPU S Y ** (Jiao et al., 2022) *** (Jiao et al., 2022) MGARD-Lambda (mga, [n. d.]) 2022 Pred. PDP+QT+LE CPU S Y - - ATC (Baert and Vannieuwenhoven, 2023) 2023 HOSVD SVD+QT+BPC+LE CPU S N ** (Baert and Vannieuwenhoven, 2023) ***** (Baert and Vannieuwenhoven, 2023) ZFP-X (Lu et al., 2023) 2023 Trans. DT+OWT+BPC+LE CPU S N **** (Lu et al., 2023) **** (Lu et al., 2023) SPERR (Li et al., 2023) 2023 Trans. DWT+QT+LE CPU S N ** (Li et al., 2023; Liu et al., 2023c) ***** (Li et al., 2023; Liu et al., 2023c) HH-NN (Huang and Hoefler, 2023) 2023 DeepL. DS+DNN+QT GPU S N * (Huang and Hoefler, 2023) ***** (Huang and Hoefler, 2023) FAZ (Liu et al., 2023c) 2023 Pred. DWT+PDP+QT+LE CPU S N ** (Liu et al., 2023c) ***** (Liu et al., 2023c) FZ-GPU (Zhang et al., 2023) 2023 Pred. QT+PDP+LE GPU S N **** (Zhang et al., 2023) ** (Zhang et al., 2023) cuSZp (Huang et al., 2023a, 2024a) 2023 Pred. QT+PDP+LE GPU S N ***** (Huang et al., 2023a, 2024a) ** (Huang et al., 2023a, 2024a) AMR-Comp (Wang et al., 2022, 2024b, 2023b, 2023a, 2024a) 2023 Pred. PDP+QT+LE CPU U N *** (Wang et al., 2024b) **** (Wang et al., 2024b) SRNN-SZ (Liu et al., 2023b) 2023 Pred. DNN+PDP+QT+LE CPU S N * (Liu et al., 2023b) ***** (Liu et al., 2023b) SZ_ADT (Lu and et al., 2023) 2023 Pred. PDP+QT+LE CPU S N *** (Lu and et al., 2023) **** (Lu and et al., 2023) HPEZ(QOZ2) (Liu et al., 2024) 2024 Pred. PDP+QT+LE CPU S N *** (Liu et al., 2024) **** (Liu et al., 2024) HAE (Le et al., 2024) 2024 DeepL. DT+DNN+QT+FTR GPU S N - **** (Le et al., 2024) SZp (szp, [n. d.]; Huang and et al., 2024a) 2024 Pred. QT+PDP+LE GPU S N ***** (Huang and et al., 2024a) ** (Huang and et al., 2024a)
@@
In this section, we describe general-purpose lossy compressors, each of which projects a specific compression pipeline composed of a series of steps or techniques/modules.
@@
Takeaway 3: We summarize the key insights based on Table 6 as follows. (1) Development of new lossy compressors evolves very quickly especially after 2021. (2) A large majority of lossy compressors (especially for the modern ones) are composed of multiple techniques. (3) CPUs are more commonly used for compressor development than GPUs and FPGAs, likely due to the stricter design requirements and more complex programming models associated with GPU and FPGA development. (4) There are far fewer compressors for unstructured data over structured data, due to the more complex processing and decorrelation required for unstructured data. (5) Compression ratio and performance form a non-trivial tradeoff. If users want to choose a compressor that meets their performance or compression ratio requirements, we highly recommend testing those rated with at least three stars in the list.
@@
SZ-family compressors. SZ (Di and Cappello, 2016; Tao et al., 2017c; Zhao et al., 2021; Liang et al., 2021a; Di and Cappello, 2018) is a prediction-based error-bounded lossy compression framework. SZ’s compression pipeline is composed of four stages: pointwise data prediction, quantization, variable-length encoding and lossless encoding. For different domain datasets and use cases, many predictors such as Lorenzo, linear regression, dynamic spline interpolation (Zhao et al., 2021), have been developed. In addition to the data prediction stage, SZ developers explored improving compression capability by other stages (e.g., different quantization methods (Liu et al., 2021b; Liu and et al., 2021; Liu et al., 2022a)). SZ adopts Huffman+Zstd (Di and Cappello, 2016; Tao et al., 2017c; Liang et al., 2018b) to compress the quantization bins, which can significantly reduce the data size (Liu et al., 2021c). SZ also supports pointwise relative error bound compression (Liang et al., 2018a), by leveraging a preprocessing step to transform the original data domain to the logarithm domain. To accelerate the throughput, an efficient fusion of logarithm was developed in (Zou et al., 2019, 2020).
@@
Based on SZ3 framework, SZ team developed many variants to adapt to diverse user-demands. For example, QoZ (quality-oriented compressor) (Liu et al., 2022b) supports compression autotuning according to user-specified quality metric targets. QoZ2.0/HPDZ (Liu et al., 2024) brings several major updates to its data prediction design, including multidimensional interpolation, interpolation re-ordering, dimension autofolding, and blockwise interpolation tuning, which can achieve 50% to 300% compression ratio improvement. FAZ (Liu et al., 2023c) combines SZ3 model with wavelet transform, also adaptively leveraging the best-fit compression techniques for each separate input and autodetermine their corresponding parameters. FAZ features extremely high compression ratios over SZ3 ( at a cost of compression throughput): it can get 3-4$\times$ compression ratio as high as that of QoZ on the SEGSalt dataset as shown in (Liu et al., 2023c). In addition, the SZ team also explored leveraging DNN to improve compression ratios under the SZ3 framework: e.g., AE-SZ (Liu et al., 2021a) leverages AutoEncoder and SRNN-SZ (Liu et al., 2023b) adopts super-resolution technique. In addition, the SZ team also developed several ultra-fast versions for both CPU and GPU devices. SZx (Yu et al., 2022), for example, composes only lightweight operations, such as bitwise operations, additions, and subtractions, so that it can get a very high compression/decompression speed (about 4$\times$ as fast as SZ2 and ZFP in general). CuSZp (Huang et al., 2023a, 2024a) can get a very high end-to-end performance on GPU for both compression and decompression (200-400GB/s on A100 in general), still maintaining a relatively high compression ratio.
@@
  • Lossless compression: compressing the shuffled bytes with a lossless compressor such as Deflate (Gzip (Deutsch, 1996)) or Zstd (Collet, 2015)).